# Gull: A Generative Multifunctional Audio Codec

### v1

@@ Abstract
We introduce Gull, a generative multifunctional audio codec. Gull is a general purpose neural audio compression and decompression model which can be applied to a wide range of tasks and applications such as real-time communication, audio super-resolution, and codec language models. The key components of Gull include (1) universal-sample-rate modeling via subband modeling schemes motivated by recent progress in audio source separation, (2) gain-shape representations motivated by traditional audio codecs, (3) improved residual vector quantization modules for simpler training, (4) elastic decoder network that enables user-defined model size and complexity during inference time, (5) built-in ability for audio super-resolution without the increase of bitrate. We compare Gull with existing traditional and neural audio codecs and show that Gull is able to achieve on par or better performance across various sample rates, bitrates and model complexities in both subjective and objective evaluation metrics. A demo page is available online [^1] https://yluo42.github.io/Gull.
@@ I Introduction
To address the aforementioned issues, we propose Gull, a generative multifunctional neural audio codec operates in the frequency domain, as the first fully-neural audio codec that supports universal sample rate, dynamic bitrate and dynamic complexity modeling of audio signals, thanks to the recent progress in band-split audio modeling for audio source separation [20, 21]. Gull splits the input spectrogram into subbands and extracts the gain-shape representations for each subband feature at each frame, and performs localized encoding via stacked residual feed forward networks (FFNs). The feature quantization process is done by an improved residual vector quantization (RVQ) module [13, 22, 15, 16] to enable faster and stabler training, introduce better reconstruction performance, and support dynamic bitrate processing. The decoder contains band-split neural network layers with dynamic width to support universal sample rate and dynamic complexity processing, and one single model can be applied to audio with any sample rate and deployed to platforms with different computational cost requirements. Adversarial training is applied to balance the signal distortion and perceptual quality of the reconstructed signal, and it also enables Gull’s built-in bandwidth extension property [23]. Experiments are designed to compare Gull with traditional audio codecs and recent neural audio codecs on both subjective and objective metrics, and we show that Gull is able to obtain on par or better performance than other systems at various signal sample rates, bitrates and model complexities with a theoretical system latency of 20 ms.
@@ II-D Quantization
We use modified subband-specific RVQ modules to quantize $\bm{\mathrm{c}}_{k,t}\in\mathbb{R}^{N}$. We use $H$ hierarchies of codebooks for each RVQ module, and the selection of codes follows the standard RVQ process, where the code with the nearest Euclidean distance to the residual embedding is selected:
$\begin{split}\bm{\mathrm{f}}^{0}_{k,t}&=\bm{\mathrm{c}}_{k,t}\\ \bm{\mathrm{e}}^{h}_{k,t}&=\bm{\mathrm{q}}^{h}_{k,j},\,j=\text{argmin}_{i}||\bm{\mathrm{q}}^{h}_{k,i}-\bm{\mathrm{f}}^{h-1}_{k,t}||_{2},\,h\geq 1\\ \bm{\mathrm{f}}^{h}_{k,t}&=\bm{\mathrm{f}}^{h-1}_{k,t}-\bm{\mathrm{e}}^{h}_{k,t}\\ \end{split}$ (3)
where $\bm{\mathrm{q}}^{h}_{k,j}\in\mathbb{R}^{N}$ denotes the $j$-th code in the codebook at hierarchy $h$ and subband $k$. We put two extra constraints on the codebooks: (1) the codes at the first hierarchy are all normalized to unit L2 norm, and (2) the codebooks at the other hierarchies always contain an all-zero code. The first constraint is equivalent to force the codebook at the first hierarchy to represent cluster centroids of an online spherical K-means clustering process [28], and the second constraint ensures that Euclidean distance between the overall quantized embedding at hierarchy $h$, i.e., $\bm{\mathrm{r}}^{h}_{k,t}=\sum_{j=1}^{h}\bm{\mathrm{e}}_{k,t}^{j}$, and the embedding to be quantized, i.e., $\bm{\mathrm{f}}^{h}_{k,t}$, is non-increasing with respect to $h$.
We normalize $\bm{\mathrm{r}}^{h}_{k,t}$ to unit L2 norm before sent to the decoder to ensure that the decoder always receives unit-norm vectors as input, and we denote the final quantized embedding tensor as $\bm{\mathrm{R}}^{h}\in\mathbb{R}^{N\times\hat{K}\times T}$. To support optional audio super-resolution property, we can set an output sample rate with its corresponding number of valid subbands $\bar{K}\geq\hat{K}$, and perform zero-padding between the $\hat{K}$-th and $\bar{K}$-th subbands to augment $\bm{\mathrm{R}}^{h}$ to shape $\mathbb{R}^{N\times\bar{K}\times T}$.
During training, we apply the exponential moving average (EMA) update [29] to the codebooks corresponding to valid input subbands, and we randomly replace codes that have not been used in 100 consecutive iterations by a random embedding in $\bm{\mathrm{F}}^{h-1}_{k}$ at the current iteration for codebook at hierarchy $h$. We also randomly select a hierarchy $h\leq H$ similar to existing works [15, 16] to allow the decoder to handle embeddings quantized at different bitrates. The quantization commitment loss is calculated at all hierarchies for the $\hat{K}$ valid subbands:
$l_{\text{commit}}=\frac{1}{\hat{K}TH}\sum_{k=1}^{\hat{K}}\sum_{t=1}^{T}\sum_{h=1}^{H}||\bm{\mathrm{c}}_{k,t}-sg[\bm{\mathrm{r}}^{h}_{k,t}]||_{2}$ (4)
@@ III-B Hyperparameters
We use slightly different hyperparameter settings for the speech model and the music model. For speech model, the operating sample rate is set to 48 kHz, and we split the input spectrogram into 10 bands, where the first 8 bands have 2 kHz bandwidth and the last 2 bands have 4 kHz bandwidth. For music model, the operating sample rate is set to 44.1 kHz, and we split the input spectrogram into 20 bands, which contain ten 400 Hz subbands, four 1 kHz subbands, four 2 kHz subbands, one 4 kHz subband, and the rest as a whole. The window and hop sizes for short-time Fourier transform (STFT) for calculating the spectrograms are 20 ms and 10 ms, respectively, which result in a theoretical system latency of 20 ms. The embedding dimension $N$ is set to 64 in both models, and the number of encoder and decoder layers is set to 4. We use LSTM layers in the decoder, and the hidden dimension is set to 128. The width of each LSTM layer (i.e., $W$) is set to 10, the dimension of FC layer in the RNN module (i.e., $N+V$ in section II-E) is set to 80 ($N=64,V=16$), and the hidden dimension in the TAC module is set to 16. For dynamic bandwidth support, we use $H=5$ hierarchies of codebooks in the RVQ modules with 12 bits in the first hierarchy and 6 bits in the rest. Table I shows the supported bitrates and actual model complexity (MACs per second) for different input and target sample rates and 3 selected decoder widths and depths. Note that no entropy coding techniques are applied to the RVQ codes, hence the bitrates reported here can be further optimized with other external modules or algorithms. For adversarial training, we use 5 discriminators with the STFT window sizes of 256, 512, 1024, 2048 and 4096, respectively, and the hop sizes are set to half of the window sizes.
@@ III-C Training objectives
The generator (i.e., codec) loss consists of a reconstruction loss, a feature matching loss, a discriminator loss, and a VQ-VAE commitment loss. The reconstruction loss operates at multiple STFT resolutions, and it includes the mean absolute error (MAE) between the complex-valued spectrograms and the mel-frequency spectrograms of the reconstructed and target signals:
$\begin{split}l_{\text{Freq}}^{w,d}&=\text{E}[\frac{|\text{Re}(\bm{\mathrm{S}}^{w,d})-\text{Re}(\bm{\mathrm{X}})|+|\text{Im}(\bm{\mathrm{S}}^{w,d})-\text{Im}(\bm{\mathrm{X}})|}{\text{mean}(|\bm{\mathrm{X}}|)}]\\ l_{\text{Mel}}^{w,d}&=\text{E}[\frac{|\text{Mel}(\bm{\mathrm{S}}^{w,d})-\text{Mel}(\bm{\mathrm{X}})|}{\text{mean}(\text{Mel}(\bm{\mathrm{X}}))}]\\ l_{\text{Rec}}&=l_{\text{Freq}}^{w,d}+l_{\text{Freq}}^{W,D}+l_{\text{Mel}}^{w,d}+l_{\text{Mel}}^{W,D}\end{split}$ (7)
@@ IV Results and analysis
Reference 96.1±0.2 94.9±0.2
MP3 6.0 65.4±1.0 24.0 79.7±1.3
Opus 6.0 58.0±1.0 24.0 93.9±0.3
EVS 7.2 92.8±0.4 24.4 94.6±0.3
Encodec 4.8 3.9G 56.1±1.0 24.0 8.1G 83.7±0.9
Gull-S 16.8 1.9G 88.4±0.8
4.8 491.5M 55.0±1.0 21.6 2.3G 91.3±0.5
24.0 2.4G 91.4±0.5
Gull-L 16.8 6.2G 92.0±0.6
4.8 1.4G 89.3±0.6 21.6 6.6G 93.9±0.3
24.0 6.7G 94.1±0.3
@@
Figure 3 shows the objective scores for various codecs on speech and music test sets. Due to the space limit and the vast amount of possible input/output sample rates and decoder model sizes of Gull, here we only select three Gull complexities, Gull-S, Gull-M and Gull-L, corresponding to decoder configurations of $\{w=1,d=1\}$, $\{w=1,d=4\}$ and $\{w=10,d=4\}$, respectively, and a few selected standard input/output sample rate configurations. Figure 2(a) and 2(b) present the evaluation of standard codec ability across different bitrates and different Gull model sizes, and Figure 2(c) and 2(d) present the evaluation of joint codec and bandwidth extension ability of different Gull input/output sample rates and model sizes.
@@
For 16 kHz speech data, we can see from Figure 2(a) that the PESQ and VISQOL scores of Gull with different model sizes are consistently better than Opus at lower bitrates ($\leq$ 8 kbps) and MP3 and Encodec at all bitrates, showing that the encoding ability of the encoder and the first few RVQ layers and the reconstruction power of the decoder are able to model the most important contents (PESQ) and maintain a reasonable sound quality (VISQOL). For middle to high bitrates, EVS performs better than all other codecs on PESQ and VISQOL, possibly due to its highly optimized codec ability to wideband speech, and Gull-L can achieve almost on par VISQOL scores as EVS. For 44.1 kHz music data, Gull configurations have worse SNR scores but much higher VISQOL scores than MP3 and Encodec. Opus performs consistently better in both measures, while EVS performs slightly worse that Gull in VISQOL.
@@
For the joint codec and bandwidth extension task for speech data, we select the input sample rate to 16 kHz and the output sample rate to 48 kHz and calculate the objective scores. For music data, we set the input sample rate to 16 or 32 kHz and the output sample rate to 44.1 kHz. The VISQOL scores for both test sets are calculated at 48 kHz sample rate. Figure 2(c) and  2(d) show that Gull can use fewer bits for transmission and leverage the power of the decoder to generate higher sample rate outputs from lower sample rate inputs, and such process can lead to improved objective scores compared to the codec-only processing scheme.
@@
Table II presents the subjective ratings of different codecs measured by MUSHRA scores. The Gull models with 16.8 kbps and 21.6 kbps for music signals are the ones with 16 kHz and 32 kHz input sample rates and bandwidth extended to 44.1 kHz output sample rate. We can see that Gull is able to achieve on-par or better performance than conventional codecs at same or lower bitrates, and an increased model complexity consistently improves the MUSHRA score. Compared to Encodec, Gull is able to achieve higher user ratings with the small model and significantly better performance with the large model, whose model size and complexity are still lower than the selected Encodec configuration.

### v2

@@ Abstract
We introduce Gull, a generative multifunctional audio codec. Gull is a general purpose neural audio compression and decompression model which can be applied to a wide range of tasks and applications such as real-time communication, audio super-resolution, and codec language models. The key components of Gull include (1) universal-sample-rate modeling via subband modeling schemes motivated by recent progress in audio source separation, (2) gain-shape representations motivated by traditional audio codecs, (3) improved residual vector quantization modules, (4) elastic decoder network that enables user-defined model size and complexity during inference time, (5) built-in ability for audio super-resolution without the increase of bitrate. We compare Gull with existing traditional and neural audio codecs and show that Gull is able to achieve on par or better performance across various sample rates, bitrates and model complexities in both subjective and objective evaluation metrics. A demo page is available online [^1] https://yluo42.github.io/Gull.
@@ I Introduction
To address the aforementioned issues, we propose Gull, a generative multifunctional neural audio codec operates in the frequency domain, as the first fully-neural audio codec that supports universal sample rate, dynamic bitrate and dynamic complexity modeling of audio signals, thanks to the recent progress in band-split audio modeling for audio source separation [20, 21]. Gull splits the input spectrogram into subbands and extracts the gain-shape representations for each subband feature at each frame, and performs localized encoding via stacked residual feed forward networks (FFNs). Motivated by the widely-used residual vector quantization (RVQ) method [13, 22, 15, 16], the feature quantization process is done by a spherical residual vector quantization (SRVQ) module where a unit-norm feature is quantized by a unit-norm codevector followed by series of rotation matrices. The decoder contains band-split neural network layers with dynamic width to support universal sample rate and dynamic complexity processing, and one single model can be applied to audio with any sample rate and deployed to platforms with different computational cost requirements. Adversarial training is applied to balance the signal distortion and perceptual quality of the reconstructed signal, and it also enables Gull’s built-in bandwidth extension property [23]. Experiments are designed to compare Gull with traditional audio codecs and recent neural audio codecs on both subjective and objective metrics, and we show that Gull is able to obtain on par or better performance than other systems at various signal sample rates, bitrates and model complexities with a theoretical system latency of 20 ms.
@@ II-D Quantization
To better quantize the unit-norm embeddings $\bm{\mathrm{c}}_{k,t}\in\mathbb{R}^{1\times N}$, we modify the standard RVQ module to spherical RVQ (SRVQ) module. For a SRVQ module with $H$ hierarchies, we set the first hierarchy to a standard VQ layer with unit-norm codebooks, and the other hierarchies to contain learnable rotation matrices defined by Householder transform. To be more specific, for the first hierarchy, the code selection process is defined as:
$\begin{split}\bm{\mathrm{e}}^{1}_{k,t}&=\bm{\mathrm{q}}^{1}_{k,j},\,j=\text{argmin}_{i}||\bm{\mathrm{q}}^{1}_{k,i}-\bm{\mathrm{c}}_{k,t}||_{2}\\ \end{split}$ (3)
where $\bm{\mathrm{q}}^{1}_{k,j}\in\mathbb{R}^{1\times N}$ denotes the $j$-th code in the codebook at subband $k$. Starting from the $h$-th hierarchy where $h\geq 2$, the code selection process is defined as:
$\begin{split}\bm{\mathrm{e}}^{h}_{k,t}&=\bm{\mathrm{e}}^{h-1}_{k,t}\bm{\mathrm{O}}^{h}_{k,j},\,j=\text{argmin}_{i}||\bm{\mathrm{e}}^{h-1}_{k,t}\bm{\mathrm{O}}^{h}_{k,j}-\bm{\mathrm{c}}_{k,t}||_{2}\\ \end{split}$ (4)
where $\bm{\mathrm{O}}^{h}_{k,j}\in\mathbb{R}^{N\times N}$ is a rotation matrix [^2] Here we treat reflection matrices as rotation matrices for similar effect. calculated by Householder transformation with unit-norm vector $\bm{\mathrm{o}}^{h}_{k,j}\in\mathbb{R}^{1\times N}$:
$\bm{\mathrm{O}}^{h}_{k,j}=\bm{\mathrm{I}}-2\bm{\mathrm{o}}^{h\top}_{k,j}\bm{\mathrm{o}}^{h}_{k,j}$ (5)
$\bm{\mathrm{o}}^{h}_{k,j}$ is a learnable parameter jointly optimized with the rest of the system. To ensure that the quantization error is non-increasing, we set $\bm{\mathrm{o}}^{h}_{k,0}=\bm{\mathrm{0}}$ to ensure that there is always an identity rotation matrix (i.e., identity mapping).
During training, we apply the exponential moving average (EMA) update [28] to the codebooks corresponding to valid input subbands at the first hierarchy, and we randomly replace codes that have not been used in 100 consecutive iterations by a random embedding in $\bm{\mathrm{c}}_{k}$ at the current iteration. We also randomly select a hierarchy $h\leq H$ similar to existing works [15, 16] to allow the decoder to handle embeddings quantized at different bitrates. For the other hierarchies, we apply the standard VQ commitment loss to update the rotation embeddings $\bm{\mathrm{o}}^{h}_{k,j}$. The overall quantization commitment loss is calculated at all hierarchies for the $\hat{K}$ valid subbands:
$l_{\text{commit}}=$ $\frac{1}{\hat{K}T}\sum_{k=1}^{\hat{K}}\sum_{t=1}^{T}||\bm{\mathrm{c}}_{k,t}-sg[\bm{\mathrm{e}}^{1}_{k,t}]||_{2}+$ (6)
$\frac{1}{\hat{K}T(H-1)}\sum_{k=1}^{\hat{K}}\sum_{t=1}^{T}\sum_{h=2}^{H}||\bm{\mathrm{c}}_{k,t}-sg[\bm{\mathrm{e}}^{h}_{k,t}]||_{2}+$ (7)
$\frac{1}{\hat{K}T(H-1)}\sum_{k=1}^{\hat{K}}\sum_{t=1}^{T}\sum_{h=2}^{H}||sg[\bm{\mathrm{c}}_{k,t}]-\bm{\mathrm{e}}^{h}_{k,t}||_{2}$ (8)
@@
We denote the final quantized embedding tensor as $\bm{\mathrm{R}}^{h}\in\mathbb{R}^{N\times\hat{K}\times T}$. To support optional audio super-resolution property, we can set an output sample rate with its corresponding number of valid subbands $\bar{K}\geq\hat{K}$, and perform zero-padding between the $\hat{K}$-th and $\bar{K}$-th subbands to augment $\bm{\mathrm{R}}^{h}$ to shape $\mathbb{R}^{N\times\bar{K}\times T}$.
@@ III-B Hyperparameters
We use slightly different hyperparameter settings for the speech model and the music model. For speech model, the operating sample rate is set to 48 kHz, and we split the input spectrogram into 10 bands, where the first 8 bands have 2 kHz bandwidth and the last 2 bands have 4 kHz bandwidth. For music model, the operating sample rate is set to 44.1 kHz, and we split the input spectrogram into 20 bands, which contain ten 400 Hz subbands, four 1 kHz subbands, four 2 kHz subbands, one 4 kHz subband, and the rest as a whole. The window and hop sizes for short-time Fourier transform (STFT) for calculating the spectrograms are 20 ms and 10 ms, respectively, which result in a theoretical system latency of 20 ms. The embedding dimension $N$ is set to 64 in both models, and the number of encoder and decoder layers is set to 4. We use LSTM layers in the decoder, and the hidden dimension is set to 128. The width of each LSTM layer (i.e., $W$) is set to 10, the dimension of FC layer in the RNN module (i.e., $N+V$ in section II-E) is set to 80 ($N=64,V=16$), and the hidden dimension in the TAC module is set to 16. For dynamic bandwidth support, we use $H=5$ hierarchies of codebooks in the SRVQ modules with 12 bits in the first hierarchy and 6 bits in the rest. Table I shows the supported bitrates and actual model complexity (MACs per second) for different input and target sample rates and 3 selected decoder widths and depths. Note that no entropy coding techniques are applied to the SRVQ codes, hence the bitrates reported here can be further optimized with other external modules or algorithms. For adversarial training, we use 5 discriminators with the STFT window sizes of 256, 512, 1024, 2048 and 4096, respectively, and the hop sizes are set to half of the window sizes.
@@ III-C Training objectives
The generator (i.e., codec) loss consists of a reconstruction loss, a feature matching loss, a discriminator loss, and a VQ-VAE commitment loss. The reconstruction loss operates at multiple STFT resolutions, and it includes the mean absolute error (MAE) between the magnitude spectrograms and the mel-frequency spectrograms of the reconstructed and target signals:
$\begin{split}l_{\text{Freq}}^{w,d}&=\text{E}[\frac{||\bm{\mathrm{S}}^{w,d}|-|\bm{\mathrm{X}}||}{\text{mean}(|\bm{\mathrm{X}}|)}]\\ l_{\text{Mel}}^{w,d}&=\text{E}[\frac{|\text{Mel}(\bm{\mathrm{S}}^{w,d})-\text{Mel}(\bm{\mathrm{X}})|}{\text{mean}(\text{Mel}(\bm{\mathrm{X}}))}]\\ l_{\text{Rec}}&=l_{\text{Freq}}^{w,d}+l_{\text{Freq}}^{W,D}+l_{\text{Mel}}^{w,d}+l_{\text{Mel}}^{W,D}\end{split}$ (11)
@@ IV Results and analysis
Reference 92.4±0.1 92.8±0.2
MP3 6.0 64.4±1.0 24.0 71.0±1.6
Opus 6.0 59.2±0.9 24.0 92.3±0.3
EVS 7.2 85.7±2.4 24.4 91.8±0.4
Encodec 4.8 3.9G 61.6±1.0 24.0 8.1G 83.7±1.0
Gull-S 2.4 315.1M 52.8±0.8 16.8 1.9G 85.8±0.9
3.6 315.1M 53.7±0.8 21.6 2.3G 89.7±0.6
4.8 491.5M 62.9±1.0 24.0 2.4G 90.2±0.5
Gull-L 2.4 1.2G 75.6±0.9 16.8 6.2G 91.7±0.4
3.6 1.2G 80.1±0.8 21.6 6.6G 90.7±0.5
4.8 1.4G 84.1±0.5 24.0 6.7G 92.5±0.3
@@
Figure 3 shows the objective scores for various codecs on speech and music test sets. Due to the space limit and the vast amount of possible input/output sample rates and decoder model sizes of Gull, here we only select three Gull complexities, Gull-S, Gull-M and Gull-L, corresponding to decoder configurations of $\{w=1,d=1\}$, $\{w=1,d=4\}$ and $\{w=10,d=4\}$, respectively, and a few selected standard input/output sample rate configurations. Figure 3(a) and 3(b) present the evaluation of standard codec ability across different bitrates and different Gull model sizes, and Figure 3(c) and 3(d) present the evaluation of joint codec and bandwidth extension ability of different Gull input/output sample rates and model sizes.
@@
For 16 kHz speech data, we can see from Figure 3(a) that the PESQ and VISQOL scores of Gull with different model sizes are consistently better than Opus at lower bitrates ($\leq$ 10 kbps) and MP3 at all bitrates, showing that the model capacity as well as the reconstruction ability of the first few SRVQ layers are able to effectively model the most important contents (PESQ) and maintain a reasonable sound quality (VISQOL). With a similar level of model complexity, Gull-L outperforms Encodec across all bitrates on PESQ and VISQOL, showing that the proposed model architecture and SRVQ module might have higher capacity on better quantizing and dequantizing the input signals. For middle to high bitrates, EVS performs better than all other codecs on PESQ, possibly due to its highly optimized codec ability to wideband speech. Gull-L can achieve almost on par VISQOL scores as EVS below 12 kbps. Among all codecs, Encodec always has the highest SNR, possibly due to the time-domain network better reconstructs the phase information.
For 44.1 kHz music data, we can see from Figure 3(b) that Gull configurations have worse SNR scores but much higher VISQOL scores than MP3 and Encodec. Opus performs consistently better than all other codecs on VISQOL with $\geq$ 20 kbps, while EVS performs slightly worse that Gull across all bitrates. Interestingly, our reproduced Encodec model performs bad on VISQOL across all bitrates, and one possible explanation might be that the bit assignment as well as the training paradigm in its RVQ module can become inefficient at higher bitrates (more codebooks), which limits its performance when the bitrate is high. As an opposite, subband-level quantization scheme in Gull allows us to flexibly find the balance between the number of subbands, number of bits in each subband and the number of RVQ levels, which bypasses the problem of assigning too many bits (i.e., codebook sizes) in one RVQ layer.
@@
For the joint codec and bandwidth extension task for speech data, we select the input sample rate to 16 kHz and the output sample rate to 48 kHz and calculate the objective scores. For music data, we set the input sample rate to 16 or 32 kHz and the output sample rate to 44.1 kHz. The VISQOL scores for both test sets are calculated at 48 kHz sample rate. Figure 3(c) and  3(d) show that Gull can use fewer bits for transmission and leverage the power of the decoder to generate higher sample rate outputs from lower sample rate inputs, and such process can lead to improved objective scores compared to the codec-only processing scheme.
@@
Table II presents the subjective ratings of different codecs measured by MUSHRA scores. The Gull models with 2.4 kbps and 3.2 kbps for speech signals are the ones with 8 kHz inputs sample rate but bandwidth extended to 16 kHz output sample rate with 1 or 2 RVQ layers, and the models with 16.8 kbps and 21.6 kbps for music signals are the ones with 16 kHz and 32 kHz input sample rates but bandwidth extended to 44.1 kHz output sample rate with 1 RVQ layer. We can see that Gull is able to achieve on-par or better performance than conventional codecs at same or lower bitrates, and an increased model complexity consistently improves the MUSHRA score. Compared to Encodec, Gull is able to achieve higher user ratings with the small model and significantly better performance with the large model, whose model size and complexity are still lower than the selected Encodec configuration.